OpenStudy (anonymous):

Medal and fan! <3 Hey guys! how do I find the equation of a vertical line passing through a point?

OpenStudy (anonymous):

My question is What is the equation of a vertical line passing through the point (−4, 7)?

OpenStudy (anonymous):

@Michele_Laino could you please help me?

OpenStudy (anonymous):

@Nnesha could you please help me? :)

Nnesha (nnesha):

 that is vertical line doesn't intersect the y-axis

Nnesha (nnesha):

vertical lines when x values are stay the same but y values are increasing so if we use slope formula to find slope \[\frac{ y_2 -y_1}{ x_2 -x_1 } =\frac{ y_2 -y_1 }{ -4-(-4) }=\frac{y_2 -y_1}{-4+4}\]i just picked two points let say we have (-4,7) then x-coordinate of 2nd line would be same ( `-4` , 5) doesn't matter what we sub for y but for vertical lines(x values would be same ) \[\frac{ y_2 -y_1}{ x_2 -x_1 } =\frac{ y_2 -y_1 }{ -4-(-4) }=\frac{y_2 -y_1}{-4+4} =\frac{y_2 -y_1}{0}\] we can't divide by 0 so slope of vertical lines would be zero

Nnesha (nnesha):

so slope of vertical lines = undefined line would be x = x-coordinate that would be ur vertical line

OpenStudy (anonymous):

Hmm so for the actual equation of the question it would just be the x-coordinate?

Nnesha (nnesha):

yes right x = x-coordinate of the given order pair :=)

OpenStudy (anonymous):

x=-4 :)?

OpenStudy (anonymous):

Thank you so much for your help :D I really appreciate it!

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha vertical lines when x values are stay the same but y values are increasing so if we use slope formula to find slope \[\frac{ y_2 -y_1}{ x_2 -x_1 } =\frac{ y_2 -y_1 }{ -4-(-4) }=\frac{y_2 -y_1}{-4+4}\]i just picked two points let say we have (-4,7) then x-coordinate of 2nd line would be same ( `-4` , 5) doesn't matter what we sub for y but for vertical lines(x values would be same ) \[\frac{ y_2 -y_1}{ x_2 -x_1 } =\frac{ y_2 -y_1 }{ -4-(-4) }=\frac{y_2 -y_1}{-4+4} =\frac{y_2 -y_1}{0}\] we can't divide by 0 so slope of vertical lines would be zero \(\color{blue}{\text{End of Quote}}\) correction `we can't divide by 0 so slope of vertical lines would be zero` i meant to say : slope of vertical lines is undefined

Nnesha (nnesha):

yes right x=-4 np :=)

OpenStudy (anonymous):

^_^ Thanks again!

Nnesha (nnesha):

o^_^o